(17 points) Match the third order linear equations with their fundamental solution sets. 1. y + 3y + 3y + y=0 2. Y - 8y +
Answer #1

So the solution is y = net A et + Bettetet tet Яз- The wronokian ent et tet -et et tet tet Y=e-t, Y2 = et tet W7,142143) ent1 = at Bl + C (again integrating y = HE The Wronnkian, 42=ti 13= 13 t til at 12:43) 372 - 6t o a -6t to i, Yz 43 are lirenaly5. y!!! O 2 AD at -> (o? ologi The Qumiliony equation in :> m(m?+1): 0 :>m=0,+ i met me 0 So the solution is 4 nt a cost of cSo the solution in y = Aest to cost te sint let y = 0 3t 92= cost, Yo - sint The wronohian W (174, 42, 43) e3b COS2 Sint 303UI y Wheren 24 +37 :+2 3ylty (13..+302 +38-+)yed euriliary equation in m3 +3m² tomt 1 = 0 dt The (m+1)3 > ma 47,-- The solutim (m-5) (m-3). c. mo 0,3,5 the solution is 50 y - At Best test Y za est Leo % = 1, Y = 0 3t The ronrakian W/4, 4₂, 43): 0.3